OpenStudy (anonymous):

factor v^2+6v+9=(v+2)(v+3)?

OpenStudy (anonymous):

no. to factor a quadratic, find factors of the last number (9) so that they add up to make the middle number (6) the correct factors for this would be (x+3)(x+3)

jimthompson5910 (jim_thompson5910):

close, v^2+6v+9=(v+3)(v+3) which condenses to v^2+6v+9=(v+3)^2

OpenStudy (anonymous):

ok how do i get the 6v?

jimthompson5910 (jim_thompson5910):

Using FOIL, we get (v+3)(v+3) v^2+3v+3v+9 v^2+6v+9

OpenStudy (anonymous):

ok then combine like terms?

jimthompson5910 (jim_thompson5910):

exactly

OpenStudy (anonymous):

ok so should i write my anwer as (v+3)(v+3) or (v+3)^2

hero (hero):

That really doesn't mater

hero (hero):

(v+3)^2 is simpler to write

OpenStudy (anonymous):

is that factored compleley?

OpenStudy (anonymous):

yes